OpenStudy (anonymous):

Georgetown and Franklin are 9.7 in. apart on a map that has a scale of 1.1in to 15 mi. How far apart are the real cities?

OpenStudy (anonymous):

proportion time.... 1.1/15 = 9.7/x (cross multiply) (1.1)(x) = (15)(9.7) 1.1x = 145.5 (divide by 1.1 on both sides to get x by itself) 1.1x/1.1 = 145.5/1.1 x = 132.27 ANSWER : Georgetown and Franklin are 132.27 miles apart :)

OpenStudy (anonymous):

thank s

OpenStudy (anonymous):

not a problem :)